### v1

# PoU: Proof-of-Use to Counter Tool-Call Hacking in DeepResearch Agents

Abstract.
Retrieval-augmented generation (RAG) agents—such as the recent DeepResearch-style systems—extend large language models (LLMs) with autonomous information-seeking capabilities via external tools. While reinforcement learning (RL) has enabled impressive multi-step reasoning performance, we reveal a previously overlooked failure mode, termed Tool-Call Hacking: agents inflate reward signals by issuing superficially correct tool calls without genuinely using the retrieved evidence. This leads to (i) mode collapse into repetitive reliance on a single source and (ii) spurious grounding, where answers are weakly supported by cited content. To address this, we propose Proof-of-Use (PoU)—an evidence-grounded RL framework that enforces verifiable causal links between retrieved evidence, reasoning traces, and final answers. PoU operationalizes this through a unified step-wise contract combining syntactic citation validation, perturbation-based sensitivity rewards, and answer–evidence alignment objectives, ensuring that tool usage remains both interpretable and functionally grounded. Across seven QA benchmarks covering in-domain, out-of-domain, and two out-of-tool-distribution tasks, PoU consistently outperforms strong DeepResearch baselines in factual accuracy, evidence faithfulness, and tool-routing balance. These findings highlight the necessity of grounding RL-trained agents not merely in task outcomes, but in the causal use of retrieved information—offering a principled path toward trustworthy retrieval-augmented reasoning.
conference: ; ; isbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Retrieval-augmented generation (RAG) agents represent a recent line of work that integrates external tools (e.g., web search)(Jin et al., 2025; Song et al., 2025; Zheng et al., 2025) into the reasoning process. By combining parametric knowledge with retrieved evidence, these agents have substantially advanced complex question answering and research assistance. However, in many knowledge-intensive settings, a single external knowledge source is often insufficient (Ma et al., 2025). This motivates scaling RAG to multi-source environments. In such settings, diverse knowledge repositories are encapsulated as callable interfaces—commonly referred to as “tools” in agent-based frameworks. Here, a “tool” denotes a retrieval interface—e.g., web search, knowledge graph, domain databases or private corpora—each with distinct coverage and reliability. The multi-source environment presents a compounded challenge: not only must agents maintain calibrated intermediate beliefs and ground their conclusions in retrieved evidence, but they must also plan strategic tool routing—selecting which tools to invoke and in what order.
Contrary to expectations, recent empirical studies in multi-source RAG show that brute-force, all-in-one retrieval can outperform prompt-based selective routing(Xu et al., 2025). This suggests that naïve prompt-based routing learned from shallow heuristics does not reliably improve the effectiveness and informativeness of retrieved content; instead, it often amplifies systemic inefficiencies and disrupts cross-source coordination, undermining the intended benefits of selectivity. In other words, when routing is weak, selectivity becomes a liability.
Building on this finding, we extend our observation to the behavioral patterns of RL-trained RAG agents. Prior works have applied reinforcement learning (RL) to enhance agents’ ability for multi‑step reasoning with retrieval tools (Zheng et al., 2025). However, we uncover a recurrent failure pattern that we term Tool‑Call Hacking—a domain‑specific form of reward hacking (Skalse et al., 2025). Specifically, the agent learns tool‑calling behaviors that boost the observed reward signal without genuinely improving grounding, which is highly susceptible to Goodhart’s Law(Majka and El-Mhamdi, 2025). We observe two consistent manifestations: (i) Mode collapse via tool overuse: policies default to a narrow subset of tools. (ii) Hallucinated tool use: tool calls are made, but the reasoning chain shows no causal dependence on the retrieved evidence. In both cases, the agent superficially satisfies the reward by producing spurious querying.
These issues highlight the need for stronger signals that foster reasoning processes deeply integrated with tool interactions in multi-source environments.
To this end, we propose an evidence-grounded agentic reinforcement learning framework, Proof-of-Use (PoU), that explicitly enforces the verifiable linkage between an agent’s reasoning process and the evidence it retrieves. Unlike conventional agentic RL paradigms that optimize for task-level success without scrutinizing how evidence is used, PoU reconstructs the entire agent–environment interaction contract into a stepwise proof schema. Within this schema, each reasoning step must first render an explicit verdict on the usefulness of retrieved evidence and then cite the specific supporting sources—thereby making the evidence–reasoning dependency a learnable and auditable interface.
This contract extends beyond intermediate reasoning: the same evidence–reasoning linkage is propagated to the answer level, where final predictions are judged by their factual consistency with the aggregated cited evidence. Moreover, the training objective itself is grounded in the contract—reward functions are computed on citation validity, perturbation sensitivity, and answer–evidence alignment, while rollout masking is selectively relaxed to preserve truly referenced snippets. Through this tightly coupled design, PoU transforms reasoning supervision from heuristic imitation into a contract-driven optimization of verifiable grounding, enabling the agent to progressively align its internal reasoning dynamics with external factual dependencies.
Scope of Tools:
In this work, we define “tools” as black-box(Ajwani et al., 2024) proxies that interface with external knowledge sources: including but not limited to open-web search, internal document collections, structured APIs, or specialized knowledge bases. This reflects a common abstraction in RAG systems, where tools are used to fetch potentially relevant information conditioned on the current query or reasoning context. While other forms of external tools (e.g., calculators, code interpreters, or simulators) have been explored in agent-based systems(Wölflein et al., 2025), our focus is on the multi-source retrieval setting where the challenge lies in tool routing and evidence integration, not symbolic execution or procedural control. This scoped definition allows us to isolate and study a critical bottleneck in retrieval-augmented reasoning: the agent’s ability to strategically call and utilize retrieval interfaces in a way that causally supports final outputs. We believe this setup is both practically relevant and theoretically rich, and we leave extension to broader tool types as future work.
Contributions.
Our work makes three key contributions:
  • We identify and formalize the tool-call hacking phenomenon of RAG agents in multi-source environments, revealing how sparse rewards and ungrounded rollouts lead to tool overuse and hallucinated grounding.
  • We propose Proof-of-Use (PoU), a contract-driven and evidence-grounded reinforcement learning framework that tightly couples the agent’s interaction protocol with causal evidence–reasoning–answer alignment, yielding verifiable and interpretable reasoning behaviors.
  • Extensive experiments across open-domain and multi-hop QA benchmarks demonstrate that PoU achieves strong performance in accuracy and faithfulness.
2. Related Works
2.1. Reinforcement Learning based Deepresearch Agent
Large language models (LLMs) are often constrained by the static knowledge embedded in their parameters, making it difficult for them to acquire the latest information in a timely manner. Consequently, when facing time-sensitive or knowledge-intensive questions, they may produce inaccurate answers or hallucinations. Building on this, RAG agents have been extended beyond static retrieval heuristics to more adaptive forms of reasoning and information seeking. Recent research has explored applying reinforcement learning (RL) to enable models autonomously plan reasoning steps, decide when to invoke what retrieval tools, and verify information across different sources (Zheng et al., 2025; Song et al., 2025). For instance, the Search-R1(Jin et al., 2025) framework allows a model to flexibly invoke search engines during answer generation, achieving a closed-loop “reason-and-search” process. The R1-Searcher(Song et al., 2025) method adopts a two-stage RL training approach that significantly enhances autonomous search capability. Similarly, ReSearch(Chen et al., 2025) enables LLMs to treat search as part of the reasoning chain even without supervised reasoning step labels. Furthermore, DeepResearcher(Zheng et al., 2025) extends the training environment to the open Internet, achieving end-to-end reinforcement learning in real-world web settings.
2.2. Reinforcement Learning Algorithms for Agent Training
In terms of optimization algorithms, mainstream approaches primarily adopt policy-gradient methods, such as Proximal Policy Optimization (PPO)(Schulman et al., 2017) and its variants. The Search-R1(Song et al., 2025) conducted a detailed comparison between PPO and the GRPO(Shao et al., 2024) in this setting. The conclusion is that GRPO converges faster, since it reduces variance by using the mean reward of multiple sampled outputs as a baseline. However, in later training stages, GRPO can exhibit reward instability or even collapse. In contrast, PPO yields a smoother and more stable training process, achieving slightly higher final performance. Ultimately, both methods reach comparable reward levels and effectively optimize Search-R1. Differently, R1-Searcher(Song et al., 2025) employs an improved version of the REINFORCE++(Hu et al., 2025) algorithm (a model-free policy gradient approach) combined with KL-divergence regularization and other stabilization techniques. In addition, the previously mentioned loss masking of retrieval results also contributes to stability—by excluding the environment’s textual feedback from gradient computation, the model avoids external noise interfering with its gradients, thus improving convergence and policy consistency. However, in our method, we tightly couple the retrieved evidence, the model’s reasoning process, and the final conclusion through an explicit contract mechanism. Under this formulation, we selectively unmask the returned evidence segments during training, thereby further encouraging the model to achieve grounded alignment between evidence and reasoning and resolve the tool-call hacking.
2.3. Reward Hacking in Reinforcement Learning
“Reward hacking” (a.k.a. specification gaming)—agents exploiting flaws in objective design to score high reward while failing the designer’s intent—was foregrounded as a core AI-safety problem in Concrete Problems in AI Safety and popularized through empirical case studies and taxonomies of gaming behaviors(Amodei et al., 2016). A formal treatment views reward hacking as arising from misspecified proxies and shows conditions under which “hackability” exists, alongside principled defenses; complementary work analyzes reward tampering, where agents manipulate the reward channel itself(Skalse et al., 2025). These phenomena are closely linked to Goodhart’s Law—over-optimization of a proxy decouples it from the true goal—and to goal misgeneralization, where systems pursue unintended objectives despite correct training feedback, both documented across deep learning and RL settings(Manheim and Garrabrant, 2019).
In post-training for LLMs, reward-model optimization introduces additional failure modes: reward model over-optimization can degrade human-judged quality as policies overfit artifacts of the learned reward. Recent studies propose constrained RLHF and composite-reward formulations to counteract these effects, while iterative RLHF empirically reduces over-optimization over successive rounds but with diminishing returns(Moskovitz et al., 2024). Process-based supervision (e.g., PRMs or step-wise CoT rewards) can provide denser, less gameable signals, though monitoring the reasoning process also creates new attack surfaces where models obfuscate or “game the monitor,” motivating hybrid outcome-plus-process objectives. Orthogonal alignment methods such as Constitutional AI reduce reliance on noisy human labels via principle-guided self-critique; while not a direct fix for reward hacking, they reframe supervision to limit exploitable gaps in reward design(Bai et al., 2022).
Modern RL-driven retrieval–reasoning agents operate in an expanded action space where the policy can call external tools (search engines, browsers, knowledge-base APIs, calculators, etc.) and thereby influence the information source used to compute downstream rewards. Therefore, we notice and formalize a class of prevalence failure modes in this expanded setting as tool-call hacking, and we aim to introduce a solution to this problem in this paper.
3. Method
A PoU agent follows an iterative reasoning–acting–observing(Yao et al., 2023) trajectory until sufficient information is accumulated to produce a final answer.
3.1. Interaction protocol and proxies
3.1.1. Initialization
Each episode begins with a user query $x$ and a predefined initial prompt that provides explicit, structured schemas for each callable tool, including their invocation formats and argument descriptions.
To simulate realistic multi-source information environments for retrieval–reasoning agents, we design and implement four categories of black-box retrieval proxies that collectively capture the diversity of external knowledge access in agentic RAG systems. Each proxy corresponds to a distinct knowledge source and interaction modality, enabling the model to reason and retrieve in a unified yet source-specific manner:
  • Web Search Proxy. This module enables the agent to access open-domain information through keyword-based retrieval. Given a textual query, it returns a structured list of webpage entries, each containing a title, URL, and snippet. For controlled evaluation, the current configuration retrieves a fixed top-$k$ set of documents per query, ensuring reproducible coverage across runs.
  • Web Browsing Proxy. This proxy performs in-depth inspection of given URLs of retrieved webpages and extracts content relevant to the current query. It returns concise, structured summaries to the reasoning model as evidence for downstream use.
  • Local Search Proxy. This proxy accesses pre-built, fixed knowledge repositories, typically used in private or professional domains such as academic papers, technical reports, enterprise data, or financial documents. Compared with web search, these repositories feature higher knowledge density and lower noise, making them particularly suitable for deeper domain-specific tasks.
  • Knowledge Graph Proxy. This proxy serves as a interface to structured, graph-based knowledge stores. Rather than employing a query-generation agent that issues executable graph queries, we adopt a semantic KG agent that abstracts subgraphs into concise, natural-language summaries. This choice prioritizes robustness and generality across heterogeneous graph schemas, allowing the proxy to be more compatible with the model’s reasoning process. The KG proxy thus functions as a lightweight retrieval utility supplying structured yet readable context when graph knowledge is required.
These tools are uniformly defined within the initialization prompt, allowing the model to understand their expected arguments and formats before the reasoning begins.
3.1.2. Proxy Response
After reasoning, the policy may issue a structured tool call.
<tool_call>JSON_CALL</tool_call>
The environment executes the call and returns a normalized proxy response with a unified schema:
<tool_response>REFERENCE_LIST_WITH_ID</tool_response>
Here, the id of each reference is automatically assigned. All four proxies adopt this schema, so the model can cite evidence via <ref>ID_1,ID_2</ref> regardless of the source.
Granularity. For the Web Search Proxy, ref items are organized at the webpage level; typically contain a page title, URL, and snippet. For the Web Browsing, Local Search, and Knowledge Graph proxies, ref items are returned at the paragraph (semantic-chunk) level to enable finer-grained grounding.
3.1.3. Reasoning Process
At the beginning of each iteration, the agent firstly performs reasoning enclosed within a <think> tag. For the first step, the model generates initial planning and reasoning solely based on the user question. From the second step onward, each <think> block must begin with a explicit verdict and citation declaration, which is treated as a unified contract:
<helpful>yes|no</helpful><ref>id1,id2,...|null</ref>
where <ref> cites the identifiers of retrieved documents from the previous tool call’s structured output.
This mechanism forces grounding each reasoning step on prior observations. From a probabilistic probing perspective, the hidden state of <helpful> token quantifies the model’s internal judgment of evidence utility, serving as a distributional anchor that aligns the model’s internal activation patterns with the conditional likelihood of evidence-relevant reasoning trajectories. By optimizing the probability mass of reasoning paths causally dependent on retrieved evidence, it suppresses spurious thoughts that diverge from verifiable context. Meanwhile, the <ref> citations anchors the evidence utility judgment to explicit external sources, establishing a probing-visible linkage between internal representations and external tool outputs. This allows each intermediate conclusion to be traceably grounded in concrete evidence.
Together, the <helpful> and <ref> mechanisms constitute a dual-channel grounding scheme that leverages their interaction as a measurable and learnable interface between probabilistic alignment and structured evidence-grounded reasoning.
When the model determines it has gathered sufficient evidence, it terminates the loop and emits:
<answer> ... </answer>
as the final response to the user.
3.2. Proof-of-use Rewards
3.2.1. Unified Step Contract for Citation Reward
According to the unified contract (denoted in 3.1.3) of each step, all proxy responses return dictionaries containing a stable id, enabling deterministic auditing. We define the Cite Reward $Reward_{cite}^{t}$ in step $t$:
(1) $\begin{split}Reward_{cite}^{t}=\begin{cases}+1,&\text{if }\texttt{parse\_ok}\land\texttt{consistency\_ok}\land\texttt{ids\_valid},\\ -1,&\text{otherwise.}\end{cases}\end{split}$
Here, $\texttt{parse\_ok}=1$ indicates syntactically valid and closed tags; $\texttt{consistency\_ok}=1$ requires $(\texttt{helpful=no}\Rightarrow\texttt{ref=null})$ and $(\texttt{helpful=yes}\Rightarrow\texttt{ref}\neq\texttt{null})$; $\texttt{ids\_valid}=1$ means all cited IDs exist in the proxy’s return list.
The rollout-level score is computed as the mean over all valid steps:
$Reward_{cite}=\begin{cases}\frac{\sum_{t=2}^{T}Reward_{cite}^{t}}{T-1},&T>1,\\[4.0pt] 0,&T=1,\end{cases}$
where $T$ is the total number of reasoning steps. This ensures that if the model answers directly without invoking any tool (T=1T{=}1), no citation reward is given.
This design separates syntax compliance from citation consistency, ensuring that every reasoning step remains both verifiable and traceable to reasoning-evidence alignment.
3.3. Reference-Linked Evidence Perturbations
To evaluate whether the model’s decision truly depends on the cited evidence rather than spurious features, we design a perturbation-based reward that measures the directional sensitivity of the model’s helpfulness prediction $p_{\theta}(\texttt{yes})$ to evidence degradation:
[…]
Specifically, for the YES case, we degrade the cited evidence set $S_{t}$ by replacing their contents with topic-unrelated snippets, and observe the probability change $\Delta q=q^{\prime}-q$, where $q=p_{\theta}(\texttt{yes}\mid R_{\mathrm{real}})$ and $q^{\prime}=p_{\theta}(\texttt{yes}\mid R_{\mathrm{pert}})$. A model that truly relies on the cited evidence should decrease its helpfulness confidence ($\Delta q<0$) after the evidence is corrupted.
For the complementary NO case, we apply the opposite perturbation: when the model initially predicts the helpfulness token as no, we randomly select one document from the proxy-returned list and replace its content with an LLM-generated snippet that is semantically relevant to the query but not necessarily factual. This synthetic “semantic lure” mimics a plausible piece of supportive evidence.
Since each rollout may contain multiple steps, the perturbation reward is computed under a fixed budget $B$:
$B=\min(T{-}1,\,B_{\max},\,\textstyle\sum_{t}\mathbb{I}[R_{cite}^{(t)}{=}1]),$
and we randomly sample $B$ steps from those with $Reward_{cite}=1$ (from step 2 to $T$) to apply perturbation. We set $B_{\max}=1$ for main experiments (thus $B=1$), and study $B=2$ in ablations to assess stability–cost trade-offs.
The final rollout-level perturbation reward is the mean of the selected $B$ results.
3.4. Answer–Citation Alignment
The large action space and delayed rewards make the agent struggle to assign credit properly, often resorting to shortcut guesses rather than grounded reasoning. To overcome the sparse credit assignment and shortcut reasoning issues, PoU directly ties the final reward to the causal consistency between the produced answer and the cited evidence through an Answer–Citation Alignment objective.
We compute the answer–evidence alignment reward as:
(3) $\begin{split}Reward_{ac}=\begin{cases}0,&\text{if }\sum_{t}\mathbb{I}[Reward_{cite}^{(t)}{=}1]=0\\ &\text{or }T{=}1,\\[3.0pt] \mathcal{J}(q,\hat{a},\mathcal{E})\in\{0,0.5,1\},&\text{otherwise.}\end{cases}\end{split}$
where $\mathcal{J}$ denotes the external LLM judge that scores whether $\hat{a}$ can be reasonably derived from the aggregated evidence set $\mathcal{E}$ of <helpful>yes</helpful> steps.
The final answer-level reward combines factual adequacy and correctness as
(4) $Reward_{a}=\frac{Reward_{ac}+Reward_{ans}}{2},$
where $Reward_{ans}$ is the F1 score between the predicted and gold answers.
We avoid overlap-based metrics (BLEU, ROUGE) for $Reward_{ac}$ since they can be trivially optimized by surface copying rather than genuine reasoning.
3.4.1. Overall Objective
The final reward integrates the three signals into a unified rollout-level objective:
(5) $Reward_{\mathrm{final}}=\begin{cases}\frac{Reward_{cite}+Reward_{pt}+Reward_{a}}{3},&\text{if format is valid,}\\ -1,&\text{otherwise.}\end{cases}$
where the overall format is considered valid only when all basic format such as <think>, <tool_call> and <answer> pass the syntactic checks.
All components are jointly optimized using Group Relative Policy Optimization (GRPO)(Shao et al., 2024). Unlike conventional agentic RL pipelines that fully mask tool outputs during training, our rollout retains the actually cited evidence snippets (those referenced via <ref>) unmasked. This aims to encourage the model to directly condition on verified causal dependency between retrieved evidence and reasoning.
3.5. Startup Fine-Tuning
To initialize a reliable agent under our PoU reasoning framework, we generate expert trajectories by a strong LLM (GPT-5). These multi-step reasoning trajectories are reformatted into our standardized interaction schema, where each turn explicitly follows the PoU contract.
For every simulated retrieval call, we emulate the output of distinct proxy tools (web_search, local_search, browser, KG_search), each returning paragraph-level evidence with stable identifiers, ensuring deterministic grounding across different information sources.
To guarantee data reliability, we apply a two-stage rejection filter to all generated trajectories. A trajectory is retained only if:
  1. (i)
    Every intermediate step passes the structural and logical validators— syntactic correctness (parse_ok), internal consistency (consistency_ok), and reference validity (ids_valid);
  2. (ii)
    The reasoning chain contains a non-trivial number of steps ($3\leq T\leq 10$), encouraging multi-hop diversity while excluding degenerate or excessively long trajectories.
The resulting dataset thus consists of verifiable, well-grounded reasoning episodes that adhere to our explicit stepwise contract. We fine-tune the base model on this curated corpus to cold-start the PoU agent.
4. Experiments
4.1. Datasets and Metrics
We evaluate PoU on a mixed in-/out-of-domain (ID/OOD) benchmark suite designed to probe open-domain QA under distributional shifts.
In-Distribution (ID)
Since our training data are drawn exclusively from HotpotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020), we report ID performance on their official development splits.
Out-of-Distribution (OOD)
To assess generalization beyond training domains, we evaluate on five additional datasets with distinct question styles and evidence structures: Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQ) (Joshi et al., 2017), MuSiQue (Trivedi et al., 2022), PopQA (Mallen et al., 2023), and Bamboogle (Press et al., 2023).
Sampling Protocol.
For comparability across datasets, we uniformly sample $512$ instances from each dev set of NQ, TQ, HotpotQA, 2Wiki, MuSiQue, and PopQA (using a fixed global seed), and include all $125$ items from the Bamboogle dev set. This yields a total of 3,1973{,}197 evaluation questions, maintaining balanced variance across datasets while avoiding confounds from dataset size. Only HotpotQA and 2Wiki are used for training; all others serve as strict OOD evaluation.
4.2. Baselines
We benchmark PoU against a diverse set of recent deep-research agents, spanning both retrieval–reasoning paradigms and RL-based search models.
(1) Retrieval–Reasoning RAG Baselines.
Routing RAG adopts a tool-routing design, where a controller dynamically selects the appropriate retriever (e.g., local corpus, web, or KG) based on query type and reasoning context. All-in-One RAG, by contrast, fuses multiple retrievers into a unified interface, jointly encoding multi-source evidence for direct reasoning without explicit routing. Both follow the ReAct (Yao et al., 2023) framework with three reasoning–retrieval turns.
(2) RL-based Search Agents.
Search-o1 (Li et al., 2025a) performs multi-hop search by iteratively generating sub-queries and retrieving concise snippets rather than full documents. Search-R1 (Jin et al., 2025) employs reinforcement learning for question answering, leveraging a Wikipedia retriever; it has both base and instruct variants. R1-Searcher (Song et al., 2025) also uses RL, but restricts retrieval to Wikipedia via site-specific filters, followed by multi-page summarization.
(3) Multi-Tool Deep-Research Agents.
DeepResearcher (Zheng et al., 2025) and Chain-of-Agent (CoA) (Li et al., 2025b) represent the latest generation of multi-source deep-research agents, operating in heterogeneous tool environments. PoU builds upon this line, introducing explicit proof-of-usefulness supervision and unified reasoning contracts to ensure verifiable evidence grounding.
4.3. Baselines
To evaluate the effectiveness of PoU, we benchmark it against a range of recent deep-research agents. Routing RAG and All-in-One RAG represent two complementary retrieval-reasoning paradigms. Routing RAG adopts a tool-routing design, where the controller dynamically decides which retriever (e.g., local corpus, web search, or knowledge graph) to invoke at each step according to the query type and reasoning context. This enables flexible retrieval planning but still relies on pre-defined routing heuristics. All-in-One RAG, in contrast, fuses all retrievers into a unified interface that jointly encodes multi-source evidence, letting the model reason directly over concatenated contexts without explicit routing. All RAG baseline following ReAct(Yao et al., 2023) with 3 turns, reflecting a fully integrated retrieval-generation pipeline.
Beyond these two RAG baselines, Search-o1(Li et al., 2025a) follows a multi-hop reasoning process where the model sequentially generates intermediate queries and retrieves short evidence snippets rather than full documents. Search-R1(Jin et al., 2025) employs reinforcement learning for question answering, using a retriever to gather Wikipedia-based evidence, with base and instruct variants depending on the initialization of the actor model. R1-Searcher(Song et al., 2025) also applies RL but confines the search scope to Wikipedia by appending site-specific constraints and summarizing multiple retrieved pages. Finally, DeepResearcher(Zheng et al., 2025) and Chain-of-Agent (CoA)(Li et al., 2025b) are two most recent deepresearch baselines with multi-source tool environments.
4.4. Implementation Details
We use GPT-4o-mini as the external judge model for evaluating answer–evidence alignment and factual consistency. All other agentic modules (browser, KG proxy, perturbations LLM) are implemented with Qwen-3-8B, deployed via vLLM inference servers. The training backbone model is Qwen-2.5-7B-Instruct. Prior to reinforcement learning, supervised fine-tuning is performed with a batch size = 256 for 3 epochs. For reinforcement learning, we adopt VERL with the GRPO algorithm. Each update uses 8 parallel rollouts, distributed across 32 × A100 GPUs. The $Reward_{pt}$ budget $B$ is 1. The maximum reasoning step is set to 10, and each retrieval call returns the top-5 documents or chunks per query. The KG proxy retrieves one-hop subgraph information from Wikidata and WikiPedia by ToG-2(Ma et al., 2025) system, providing semantically condensed neighborhood summaries as context. In the OODT experiments, we additionally introduce a local biochemical Proxy, which is built from a biochemical corpus derived from PubMed and serves as a local retrieval interface, serving as the fixed local knowledge base for domain-specific retrieval. The local_proxy is turn-off for non-medical domain tasks.
4.5. Main Results
Model Env. HotpotQA 2Wiki
F1 LM F1 LM
RAG
Routing RAG Multi-source 30.8 36.0 14.4 18.2
All-in-one RAG Multi-source 34.5 40.2 23.2 26.4
Search-o1* Local 31.6 40.8 28.6 32.8
Search-o1 Web Search 33.0 42.4 30.9 37.7
RL Training RAG
Search-R1-base Local 55.9 63.0 44.6 47.9
Search-R1-instruct Local 45.7 52.5 43.4 48.8
R1-Searcher Web Search 44.8 53.1 59.4 65.8
CoA Multi-source 43.9 50.2
DeepResearcher Web Search 52.8 64.3 59.7 66.6
Ours
PoU Multi-source 53.0 65.4 67.8 76.8
Table 1. In-domain evaluation on HotpotQA and 2Wiki datasets. All results are reported with F1 and LLM-based evaluation (LM) metrics. Bold numbers indicate the best performance in each column. Underlined numbers indicate the second best.
Method Env. NQ TQ MuSiQue Bamboogle PopQA
F1 LM F1 LM F1 LM F1 LM F1 LM
RAG
Routing RAG Multi-source 34.8 42.9 51.0 59.6 6.5 8.2 17.0 17.4 35.1 39.9
All-in-one RAG Multi-source 37.5 48.3 60.2 68.8 9.7 11.8 23.5 25.1 46.9 48.8
Search-o1* Local 34.5 57.4 52.6 61.1 16.8 21.3 35.8 38.4 36.9 42.4
Search-o1 Web Search 32.4 55.1 58.9 69.5 14.7 19.7 46.6 53.6 38.3 43.4
RL Training RAG
Search-R1-base Local 45.4 60.0 71.9 76.2 26.7 27.5 56.5 57.6 43.2 47.0
Search-R1-instruct Local 33.1 49.6 44.7 49.2 26.5 28.3 45.0 47.2 43.0 44.5
R1-Searcher Web Search 35.4 52.3 73.1 79.1 22.8 25.6 64.8 65.6 42.7 43.4
CoA Multi-source 43.9 63.3 22.3 49.6 46.5
DeepResearcher Web Search 39.6 61.9 78.4 85.0 27.1 29.3 71.0 72.8 48.5 52.7
Ours
PoU Multi-source 45.5 67.2 77.6 89.6 29.0 34.1 69.5 74.6 55.7 61.2
Table 2. Performance comparison on five out-of-domain QA benchmarks: NQ, TQ, MuSiQue, Bamboogle, and PopQA. All results are reported with F1 and LLM-based evaluation (LM) metrics. Bold numbers indicate the best performance in each column. Underlined numbers indicate the second best.
Across all seven datasets—including both in-domain (HotpotQA, 2Wiki) and out-of-domain (NQ, TQ, MuSiQue, Bamboogle, PopQA) benchmarks—our proposed Proof-of-Usefulness (PoU) agent consistently outperforms prior RAG and RL-trained baselines. This advantage is not limited to the in-domain subset: on the out-of-domain benchmarks, PoU achieves clear gains in both factual F1 and LLM-as-judge (LM) metrics, demonstrating superior generalization beyond its training distribution.
Interestingly, even on datasets such as NQ and TQ—which serve as in-domain data for the DeepResearcher baseline—our model achieves higher overall scores. This observation suggests that PoU’s improvement does not rely on memorizing domain-specific knowledge, but rather stems from stronger evidence–reasoning alignment. The model effectively learns to utilize retrieved evidence and to generalize its reasoning strategies to novel domains, indicating that PoU captures transferable reasoning competence rather than dataset-specific experience.
We also observe that All-in-One RAG consistently outperforms Routing RAG across all evaluation sets. This finding corroborates the observation in RAG in the Wild that naïve or heuristic routing—without explicit optimization or supervision—tends to be inefficient or even counterproductive. Although routing appears intuitively beneficial by narrowing the retrieval scope and reducing search cost, it also risks prematurely discarding relevant information.
In our context, the PoU agent operates in a multi-source environment where routing becomes an additional latent variable in the decision space. Compared with traditional deep-research agents, this introduces a new layer of complexity: each reasoning step must now decide not only what to retrieve but also where to retrieve from. Such an expanded action space grows exponentially with the number of information sources, leading to inefficient exploration when the routing policy is not explicitly optimized—precisely the phenomenon we aim to address through structured reasoning and contract-based training under the PoU framework. By explicitly encoding citation validity and reasoning consistency in every decision step, PoU regularizes the search process and prevents the agent from diverging in the enlarged action space. This design enables the agent to leverage multi-source evidence effectively without suffering from the instability and inefficiency commonly observed in naive routing-based systems, ultimately yielding superior performance and generalization ability in multi-source environments.
4.6. Ablation Study
4.6.1. Rewards Ablation
Figure 1 presents the ablation study of PoU under different reward configurations, where each reward component is selectively removed or adjusted. We observe that increasing the perturbation budget to B=2B{=}2 leads to a more stable training trajectory and a higher final performance. This confirms the effectiveness of the perturbation-based reward $Reward_{pt}$: by actively perturbing the cited evidence, the model learns to rely on genuinely supportive information rather than superficially acknowledging tool responses. Such perturbation helps prevent the agent from “pretending” to use tool feedback, thereby maintaining a healthy and optimizable training regime instead of exploiting spurious reward shortcuts. However, since the rollout cost of GRPO scales approximately multiplicatively with the perturbation budget, we adopt B=1B{=}1 as a balanced trade-off between computational efficiency and stability.
Furthermore, removing the answer–citation alignment term $Reward_{ac}$ results in unstable training and a sudden collapse in later stages. This indicates that linking the final answer with intermediate evidence provides an essential smoothing and grounding signal for the policy. Without this alignment, the model easily drifts away from verifiable reasoning behaviors, demonstrating the critical role of answer–evidence coupling in maintaining robust optimization.
Figure 1. Ablation study of PoU. Comparison under different reward configurations: default perturbation budget B=1B{=}1, extended B=2B{=}2, and variants removing $Reward_{pt}$ or $Reward_{ac}$. Increasing the perturbation budget to B=2B{=}2 yields more stable convergence and a higher upper bound, validating the effectiveness of the perturbation-based reward $Reward_{pt}$. Perturbation prevents the agent from “pretending” to read tool feedback and keeps optimization in a healthy, learnable regime rather than exploiting spurious shortcuts. However, since GRPO rollout cost scales multiplicatively with $B$, we adopt B=1B{=}1 as a practical trade-off. Removing the answer–citation alignment term $Reward_{ac}$ causes the model to collapse in later stages, confirming that explicit evidence–answer alignment provides a crucial smoothing and grounding signal for stable training.
4.6.2. Out-of-domain (OODT) Experiments
To further validate the generalization ability of our approach in vary environments, we conducted additional experiments under an out-of-domain (OOD+OODT) setting. Specifically, we introduced a fourth proxy—a biomedical local search proxy—which was never seen during training. We then compared our model with existing baselines on two biomedical QA datasets, BioASQ(Krithara et al., 2023) and PQArefEval(Bašaragin et al., 2024). This setup constitutes an OOD+OODT scenario (out-of-domain data and out-of-domain tools), enabling us to assess how well the model generalizes when both the data distribution and the available tool set differ from those encountered during training.
Dataset Zero-shot RAG DeepResearcher+ DeepResearcher PoU (Ours)
BioASQ 13.5 25.8 37.1 38.9 45.6
PQArefEval 18.4 31.5 42.4 42.1 48.7
Table 3. OOD+OODT evaluation on biomedical QA datasets (BioASQ, PQArefEval). The experiments introduce a fourth proxy, a biomedical local search proxy, to test generalization across both domain and toolset shifts.
As shown in Table 3, even though DeepResearcher+ was trained in a more complex 3-tool environment, its performance is notably inferior to its original 2-tool counterpart (DeepResearcher) on both biomedical datasets. In contrast, our proposed PoU model achieves the best results by a clear margin, demonstrating strong domain-level and tool-level generalization. This finding highlights an important phenomenon: as the number of available tools increases, the action space of the agent expands combinatorially. Without explicit process-level reward signals, such enlarged search space can easily lead to policy collapse or unstable reasoning behavior. In contrast, the PoU training paradigm—through explicit evidence-reasoning alignment and structured reward design—successfully stabilizes the learning process and preserves robustness under tool and domain distribution shifts, indicating the Tool-call Hacking is significantly mitigated.
4.6.3. Tool-Call Behavior Analysis
To assess how expanding the tool set changes agent behavior, we trained Deepresearcher+ with one more tool (Knowledge Graph) from scratch under the same data and hyperparameters as the original Deepresearcher, then uniformly sampled 1,0001{,}000 queries from seven test sets and counted tool-call usage across three categories: Web Search (Web), Knowledge Graph (KG), and Web Browser (Browser). Table 4 summarizes the call ratios.
Table 4. Tool-call ratio (%) over 1,0001{,}000 sampled queries from seven test sets. Columns are Web Search (Web), Knowledge Graph(KG), and Web Browser (Browser). Deepresearcher has no KG tool (0.0).
Model Web KG Browser
Deepresearcher 88.4 - 11.6
Deepresearcher+ 94.6 2.2 3.2
PoU 51.8 27.5 20.7
The results in Table 4 reveal distinct behavioral patterns across models. When trained with only the format and answer rewards (as in the original Deepresearcher setup), the agent tends to overuse available tools once the tool set expands. After adding the Knowledge Graph interface, Deepresearcher+ exhibits a pronounced bias toward redundant Web Search calls (94.6%) and a near disappearance of Web Browser usage (3.2%). This behavior indicates a collapse of tool specialization: although the browser is logically designed to serve as a follow-up step for Web Search—retrieving detailed content from previously retrieved URLs—the model largely ignores this expected hierarchy and instead repeatedly queries the Web Search proxy. Such uncontrolled escalation of retrieval actions suggests severe Tool-call Hacking when training lacks explicit evidence-usage constraints.
In contrast, the PoU-trained agent demonstrates a markedly more balanced tool distribution. Each proxy—Web Search, Knowledge Graph, and Browser—maintains a complementary role (51.8%, 27.5%, 20.7%, respectively), implying that the model actively schedules heterogeneous retrieval utilities based on their strengths: the Web Search proxy for broad discovery, the Knowledge Graph for exploratory search, and the Browser for fine-grained verification. This balanced behavior highlights that the PoU reward design effectively regularizes tool usage, encouraging agents to retrieve, reason, and cite evidence in a coordinated manner rather than relying on repetitive search actions.
4.6.4. Extreme Ablation
To further examine robustness and causal grounding, we replaced every <tool_response> with the token “content”, removing all semantics while keeping structure.
Model Behavior under ”content” responses
PoU <think>”No evidence; Try other tools.” $\rightarrow$ <helpful>no</helpful> $\rightarrow$ <ref>null</ref> $\rightarrow$ <tool_call> KGSearch(q)(q).
$\Rightarrow$ Detects missing evidence and switches tools/parameters.
[…]
$\Rightarrow$ Hallucinates evidence and continues fictitious reasoning.
Table 5. Extreme ablation test where all tool responses are replaced by “content”. Only PoU detects missing evidence and adapts its strategy.
5. Conclusion
In this work we identified tool‑call hacking as a pervasive failure mode in multi‑source retrieval‑augmented agents and introduced Proof‑of‑Use (PoU)—a contract‑driven, evidence‑grounded RL framework that enforces step‑wise alignment between retrieved evidence, intermediate reasoning, and final answers. In PoU this contract contains explicit helpfulness/citation declarations, a citation‑format reward, reference‑linked perturbation sensitivity, and an answer–citation alignment objective, coupled with selective unmasking of actually cited snippets during training. Across diverse in‑domain and out‑of‑domain QA benchmarks, PoU yields consistent gains in both F1 and LLM‑as‑judge scores over strong deep‑research baselines, while exhibiting balanced and purposeful tool routing and robustness under domain even tools shifts. Limitations include reliance on an external LLM judge for alignment scoring, and a current focus on retrieval tools rather than procedural tools. By the ablation studies, we clearly demonstrate the rationality and necessity of each reward design. Future work will explore learned or weakly supervised judges, adaptive perturbation curricula, cost‑ and latency‑aware routing, extensions to calculators/coders/simulators, and robustness against adversarial or privacy‑sensitive tool outputs. Overall, PoU offers a simple and auditable path toward trustworthy multi‑source reasoning by making evidence use not only observable but optimizable.
Ethical Use of Data and Informed Consent
This study does not involve human participants or personally identifiable information. All datasets used (e.g., HotpotQA, PubMed, and other publicly available corpora) are released under open research licenses and comply with their respective terms of use. All web content accessed complied with robots.txt and site terms, and was used solely for non-commercial research. No additional ethical approval was required.